# DESI-Like Hubble Expansion From Staged Symmetry Breaking

### v1

@@ I Introduction
In this work, we build on the idea by providing an example of a renormalizable model that can exhibit similar behavior. While it may at first seem that one could construct any arbitrary variation of the equation of state parameter by reverse-engineering a potential like the one given above, this is not the case. For a model to be physically sound as a quantum field theory, it needs a UV completion. It is desirable for a theory to emerge from a natural, renormalizable Lagrangian. One can treat some non-renormalizable theories as effective field theories, though without a UV completion, such a model is potentially unphysical. It is desirable that an effective field theory emerges naturally as the low-energy limit of a renormalizable theory, as non-renormalizable theories produce divergences that cannot be removed using a finite number of counterterms. We find that the space of possible models is heavily constrained by the desire to avoid unphysical or problematic constructions, and we identify at least one model that shows some possible promise.
@@ II Discussion of Possible Models
We wish to produce a dark matter model whose equation of state parameter shifts between four values: $w_{1},w_{2},w_{3},$ and $w_{4}$. To do this, we consider a decay cascade of the form $\rm{A}\to\rm{B}\to\rm{C}\to\rm{D}$. We consider a cascade rather than a complicated potential with one field, because renormalizability requires that we have no operator with a mass-dimension greater than four, which forbids terms like $\lambda\phi^{6}$. This cascade concept is similar to the dynamical dark matter considered by Dienes [50, 51, 52], though our work contains a finite number of stages, rather than an infinite number, and is motivated by DESI observations. In addition to our fiducial model, we briefly discuss alternatives and outline the problems with them.
@@ II.2 Four Complex Scalar Fields With Global Symmetry
We can forbid unwanted decays like $\rm{A}\to 2\rm{C}$ by assigning appropriate charges to the fields. This can be accomplished by upgrading the real scalar fields to complex scalar fields with a global $\rm{U}(1)$ symmetry. This can be used to forbid the unwanted one-loop decays and avoid the naturalness problem of omitting terms like $g_{\rm{AD}}\rm{AD}^{2}$, as charge assignments forbid them.
Such a model still has a couple issues, however. Global symmetries are believed to be inconsistent once embedded in a full theory of quantum gravity [53]. Even without appealing to quantum gravity consistency, one can argue that gauge symmetries are natural. Global symmetries expected to be only approximate at best, with gauge symmetries taking primacy [54]. This conjecture can be motivated by the AdS/CFT correspondence [53], which can be regarded as a toy model for quantum gravity where anti-de Sitter space (AdS) is dual to a conformal field theory (CFT). Our Universe is not anti-de Sitter, as the cosmological constant is believed to be positive, rather than negative, though the arguments used for AdS/CFT could perhaps be generalized to flat-space holography, which would describe a Universe with a vanishing cosmological constant, and would be a more useful approximate model for our Universe.
When considering continuous global symmetries, the case that they are inconsistent can also be motivated by the black hole information paradox, in which case the argument can be generalized beyond the AdS/CFT correspondence [55]. One can loosely motivate this by arguing that continuous global symmetries imply conserved charges that are destroyed by black hole evaporation, therefore continuous global symmetries should not exist in nature, as we expect black hole evaporation to be unitary. This argument is more restrictive than the AdS/CFT result, as it applies to continuous global symmetries, such as U(1), but not global symmetries in general. More importantly, the theory will suffer the same problem as the model with four real scalar fields in that it permits a mixture of A, B, C, and D fields, and that suppresses the effect on H(z). One could consider spontaneously breaking a global $\rm{U}(1)\times\rm{U}(1)\times\rm{U}(1)$ symmetry to sequentially turn on two-body decays of the form $\rm{A}\to 2\rm{B}$, though this would still incorporate an undesirable global symmetry.
@@ II.3 Our Fiducial Model
We have four complex scalar fields A, B, C, and D, with four dark Higgs fields $\rm{H}_{1}$, $\rm{H}_{2}$, $\rm{H}_{3}$, and $\rm{H}_{4}$. We use complex scalar fields because charge conservation allows us to enforce a decay chain, and because complex scalars enable a renormalizable path to a temporarily negative equation of state parameter with an analytic potential, whereas real scalar fields would require non-analytic terms, such as $|\phi|$, or a field that slow-rolls, which would act as quintessence, rather than dark matter. We note that potentials involving $|\phi|$ have been studied within the Wilsonian framework, though they remain problematic for perturbation theory [56]. One might initially suspect that a scalar field will not clump as dark matter because the kinetic terms could imply a sound speed of $c$, though there is no such problem [18], which we discuss further in the Appendix.
Since the theory has a U(1) $\times$ U(1) $\times$ U(1) $\times$ U(1) gauge group, we must introduce four gauge bosons to allow for gauge invariance: $\alpha_{\mu}$, $\beta_{\mu}$, $\delta_{\mu}$, and $\gamma_{\mu}$, where we define the Faraday tensors as $\Psi_{\mu\nu}=\partial_{\mu}\Psi_{\nu}-\partial_{\nu}\Psi_{\mu}$, and the covariant derivative as $\mathcal{D}_{\mu}\equiv\partial_{\mu}-i(g_{1}q_{1}\alpha_{\mu}+g_{2}q_{2}\beta_{\mu}+g_{3}q_{3}\gamma_{\mu}+g_{4}q_{4}\delta_{\mu})$. We note that similar gauge groups can naturally arise in string theory [57]. We also note that anomalies will not be present, as the model does not incorporate fermions [58]. We can express $\mathcal{L}$ as follows, where we have written the linear “tadpole” term for C in boldface to distinguish it for later discussions. For simplicity, we assume any mixed quartics of the form $(\mathrm{H}_{1}^{\dagger}\mathrm{H}_{1})(\mathrm{H}_{2}^{\dagger}\mathrm{H}_{2})$ have small coupling constants, and can be neglected.
$\mathcal{L}\supset$ $\;\mathcal{D}^{\mu}\mathrm{A}^{\dagger}\mathcal{D}_{\mu}\mathrm{A}-m_{\mathrm{A}}^{2}\,\mathrm{A}^{\dagger}\mathrm{A}+\mathcal{D}^{\mu}\mathrm{B}^{\dagger}\mathcal{D}_{\mu}\mathrm{B}-m_{\mathrm{B}}^{2}\,\mathrm{B}^{\dagger}\mathrm{B}-\lambda_{\mathrm{B}}(\mathrm{B}^{\dagger}\mathrm{B})^{2}$ (4)
$+\mathcal{D}^{\mu}\mathrm{C}^{\dagger}\mathcal{D}_{\mu}\mathrm{C}-m_{\mathrm{C}}^{2}\,\mathrm{C}^{\dagger}\mathrm{C}+\mathcal{D}^{\mu}\mathrm{D}^{\dagger}\mathcal{D}_{\mu}\mathrm{D}-m_{\mathrm{D}}^{2}\,\mathrm{D}^{\dagger}\mathrm{D}-\lambda_{\mathrm{D}}(\mathrm{D}^{\dagger}\mathrm{D})^{2}$
$-\Big(\bm{\kappa\,\mathrm{C}\,\mathrm{H}_{1}^{\dagger}}+g_{\mathrm{AB}}\mathrm{A}\mathrm{B}^{2}\mathrm{H}_{2}^{\dagger}+g_{\mathrm{BC}}\mathrm{B}\mathrm{C}^{2}\mathrm{H}_{3}^{\dagger}+g_{\mathrm{CD}}\mathrm{C}\mathrm{D}^{2}\mathrm{H}_{4}^{\dagger}+\text{h.c.}\Big)$
$-\sum_{\Psi_{\mu}\in\{\alpha_{\mu}\beta_{\mu},\gamma_{\mu},\delta_{\mu}\}}[\tfrac{1}{4}\!\Psi_{\mu\nu}\Psi^{\mu\nu}]+\sum_{i=1}^{4}\!\Big[\mathcal{D}^{\mu}\mathrm{H}_{i}^{\dagger}\mathcal{D}_{\mu}\mathrm{H}_{i}+m_{\mathrm{H}_{i}}^{2}\,\mathrm{H}_{i}^{\dagger}\mathrm{H}_{i}-\lambda_{\mathrm{H}_{i}}(\mathrm{H}_{i}^{\dagger}\mathrm{H}_{i})^{2}\Big]$
@@
Each time a U(1) is spontaneously broken, a dark Higgs field gains a non-zero vacuum expectation value. This then sequentially turns on the terms of the form $\rm{C}$, $\rm{A}\rm{B}^{2}$, $\rm{B}\rm{C}^{2}$, and $\rm{C}\rm{D}^{2}$. We note that the mass of the daughter must be very close to half the mass of the parent, both to ensure that decay products are cold, and to suppress three-body decays involving a massless dark Higgs before symmetry breaking. This necessity is one of the greatest obstacles we face in providing a natural model whose behavior resembles that desired by Chen and Loeb [18].
The terms $\rm{A}\rm{B}^{2}$, $\rm{B}\rm{C}^{2}$, and $\rm{C}\rm{D}^{2}$ enable decays from one field to another, whereas the linear term (see boldface term in the action) allows C to have a negative equation of state parameter. The equation below yields the equation of state parameter $w(R)$ for a complex scalar field, which can be written in terms of its magnitude, $R$, and phase factor, $\phi$, giving $Re^{i\phi}$ [59].
@@
Here, the equation of state parameter is $w=1/3$ for a quartic potential, $w=0$ for a quadratic potential, and w=1/3w=-1/3 for a linear potential. If the field C is oscillating in a potential that has both a linear term and a quadratic mass term, but where the linear term dominates, one obtains $w\approx-1/3$. A linear potential for a complex scalar is then analogous to a potential $|\phi|$ for a real scalar $\phi$. For a complex scalar, $R$ is strictly positive, so a linear term ensures that the field can oscillate in the potential; this would not be the case if the scalar were real, because a real scalar requires $|\phi|$ instead of $\phi$. A potential of the form $|\phi|$ is undesirable because it is non-analytic.
Note that the linear term for C turns on before any of the decay channels, as the associated U(1) is broken first, though this term has no effect on the system until C is produced from decays later on. At first, one may think that the linear term could be removed by shifting the field C, though this is not possible. To illustrate what we mean by a “shift”, consider the following terms, where we use a real scalar field for simplicity. We see that the constant on the right-hand side can be absorbed into the vacuum energy, so the linear term can be removed via a field redefinition.
$\frac{1}{2}m_{\mathrm{C}}^{2}\mathrm{C}^{2}+\kappa\mathrm{C}=\frac{1}{2}m_{\mathrm{C}}^{2}\bigg(\mathrm{C}+\frac{\kappa}{m_{\mathrm{C}}^{2}}\bigg)^{2}-\frac{\kappa^{2}}{2m_{\mathrm{C}}^{2}}$ (6)
For our decay chain, this is not possible, as such a shift in C would introduce new couplings into the action, thus changing the physics. The first reason for this is that the linear term involves a coupling to the dark Higgs, $\mathrm{H}_{1}$. A shift of the form $\kappa(\mathrm{C}+\epsilon)\mathrm{H}_{1}^{\dagger}$ thus effectively generates a new term of the form $\epsilon\mathrm{H}_{1}^{\dagger}$. The second reason can be seen in the decay terms. A shift there yields $\mathrm{B}(\mathrm{C}+\epsilon)^{2}=\mathrm{B}\mathrm{C}^{2}+2\epsilon\mathrm{B}\mathrm{C}+\epsilon^{2}\mathrm{B}$, so the shift generates a new interaction term of the form $\mathrm{B}\mathrm{C}$, as well as a term linear in B.
One can also note that the linear term has an effect that can be roughly thought of as resembling a step-like shift in the effective vacuum energy. To see this, consider a massive scalar B that decays to a massive scalar C, where only C has a linear term. Both fields have parabolic potentials, though the minima of the potentials are at different places. (To illustrate this, we plot the functions $f(x)=\frac{1}{2}x^{2}$ and $g(x)=\frac{1}{2}x^{2}+x$ in Figure 1.) If one had a scalar field in isolation, whose potential had both a linear term and a mass term, then one could shift the field, absorb a resulting constant term into the vacuum energy, and end up with a purely quadratic potential, though this does not work in a decay chain. The linear term thus produces physical effects.
Figure 1: We plot two parabolic functions, $f(x)=\frac{1}{2}x^{2}$ and $g(x)=\frac{1}{2}x^{2}+x$, to illustrate that adding a linear term to a potential shifts its minimum.
@@ II.4 Self-Interacting Dark Matter Candidate From Our Fiducial Model
After the first symmetry breaking yields a dark Higgs with a mass of $\mathcal{O}$(MeV), the term $\mathcal{D}^{\mu}\rm{H_{1}}^{\dagger}\mathcal{D}_{\mu}\rm{H_{1}}$ produces an interaction of the form $gh\alpha_{\mu}\alpha^{\mu}$, which permits an $\alpha\alpha\to\alpha\alpha$ scattering processes via a scalar mediator, as shown below. Note that $h$ is the radial mode of the dark Higgs, after symmetry breaking gives it a mass. For $\alpha\sim\rm{GeV}$ and $h\sim\rm{MeV}$, our fiducial model produces a natural SIDM candidate, one that could conceivably address some of the issues with CDM if $\alpha$ comprises the bulk of the dark matter. We show the $t$ and $u$-channel Feynman diagrams in Figure 2.
@@ III Methods
We have equation of state parameters $w_{\rm{A}}=0$, $w_{\rm{B}}=\frac{1}{3}$, $w_{\rm{C}}=-\frac{1}{3}$, and $w_{\rm{D}}=\frac{1}{3}$. We activate decays at redshifts of 3, 2, and 0.4. We use $\Gamma_{\rm{A}}=\Gamma_{\rm{B}}=\Gamma_{\rm{C}}=10^{-15}$ $\rm{sec^{-1}}$. (See the Appendix for the equation that expresses the decay rate in terms of the masses and coupling constants.) Note that for simplicity, we neglect any three-body decays. Future work could incorporate this. In the equations below, $\rho_{\mathrm{A}}$, $\rho_{\mathrm{B}}$, $\rho_{\mathrm{C}}$, and $\rho_{\mathrm{D}}$ are the densities of the complex scalars in the cascade, whereas $\rho_{\mathrm{O}}$ is the density of the other matter (baryons+other dark matter), $\rho_{r}$ is the radiation density, and $\rho_{\Lambda}$ is the dark energy density.
$\dot{\rho}_{\rm A}+3H(1\!+\!w_{\rm A})\rho_{\rm A}$ $=-\Theta(t\!-\!t_{1})\Gamma_{\rm A}\rho_{\rm A}$ (7a)
$\dot{\rho}_{\rm B}+3H(1\!+\!w_{\rm B})\rho_{\rm B}$ $=\Theta(t\!-\!t_{1})\Gamma_{\rm A}\rho_{\rm A}$
$\quad-\Theta(t\!-\!t_{2})\Gamma_{\rm B}\rho_{\rm B}$ (7b)
$\dot{\rho}_{\rm C}+3H(1\!+\!w_{\rm C})\rho_{\rm C}$ $=\Theta(t\!-\!t_{2})\Gamma_{\rm B}\rho_{\rm B}$
$\quad-\Theta(t\!-\!t_{3})\Gamma_{\rm C}\rho_{\rm C}$ (7c)
$\dot{\rho}_{\rm D}+3H(1\!+\!w_{\rm D})\rho_{\rm D}$ $=\Theta(t\!-\!t_{3})\Gamma_{\rm C}\rho_{\rm C}$ (7d)
@@
We integrate the system of equations from recombination at $z_{\rm rec}=1090$ to $z=0$, where we specify our initial conditions by setting $h_{i}$ = 0.674, where $\rm{H_{0}}$ is $100\rm{\frac{km}{secMpc}}$$h_{i}$ in units of $\rm{sec}^{-1}$. The radiation $\Omega_{r}$, matter $\Omega_{m}$, and $\Omega_{\Lambda}$ components are dictated by the equations written below [60, 61]:
@@
$\rho_{\Lambda,i}=\Omega_{\Lambda}\rho_{\rm crit,0},$ (15)
III.1 The Hubble Parameter Evolution for our Fiducial Model
We use the equation below to calculate H(z) for a Universe with cold dark matter and dynamical dark energy [62]. We plot H(z) from our model alongside $w_{0}w_{a}$CDM, which has the equation of state parameter $w_{\rm{DE}}(a)=w_{0}+w_{a}(1-a)$ for dark energy, to show that our model could display qualitative consistency with the DESI DR2 trends. Similar to Chen and Loeb [18], who deferred a fit to DESI data to a separate study [19], we leave a full parameter fit to DESI results for future work. Such a fit is beyond the scope of the current study to explore the space of potentially viable models.

### v2

@@ I Introduction
In this work, we build on the idea by providing an example of a more natural effective field theory that can exhibit similar behavior. While it may at first seem that one could construct any arbitrary variation of the equation of state parameter by reverse-engineering a potential like the one given above, this is not the case. For a model to be physically sound as a quantum field theory, it needs a UV completion. It is desirable for a theory to emerge from a natural, renormalizable Lagrangian. One can treat some non-renormalizable theories as effective field theories, though without a UV completion, such a model is potentially unphysical. We are unable to produce a field theory that exhibits the desired behavior while being fully renormalizable, though in our case, all interactions are renormalizable, except one.
@@ II Discussion of Possible Models
We wish to produce a dark matter model whose equation of state parameter shifts between four values: $w_{1},w_{2},w_{3},$ and $w_{4}$. To do this, we consider a decay cascade of the form $\rm{A}\to\rm{B}\to\rm{C}\to\rm{D}$. This cascade concept is similar to the dynamical dark matter considered by Dienes [51, 52, 53], though our work contains a finite number of stages, rather than an infinite number, and is motivated by DESI observations. In addition to our fiducial model, we briefly discuss alternatives and outline the problems with them.
@@ II.2 Four Complex Scalar Fields With Global Symmetry
We can ensure a clean shift of energy density from one field to another (so the effect on H(z) is not suppressed) by using complex scalars with appropriate charges. One could spontaneously break a global $\rm{U}(1)\times\rm{U}(1)\times\rm{U}(1)$ symmetry to sequentially turn on two-body decays of the form $\rm{A}\to\mathrm{B}^{\dagger}\mathrm{B}$. Sequentially turning on decays in this way helps to ensure a clean shift of energy density between fields.
Such a model still has an issue, however. Global symmetries are believed to be inconsistent once embedded in a full theory of quantum gravity [54]. Even without appealing to quantum gravity consistency, one can argue that gauge symmetries are natural. Global symmetries expected to be only approximate at best, with gauge symmetries taking primacy [55]. This conjecture can be motivated by the AdS/CFT correspondence [54], which can be regarded as a toy model for quantum gravity where anti-de Sitter space (AdS) is dual to a conformal field theory (CFT). Our Universe is not anti-de Sitter, as the cosmological constant is believed to be positive, rather than negative, though the arguments used for AdS/CFT could perhaps be generalized to flat-space holography, which would describe a Universe with a vanishing cosmological constant, and would be a more useful approximate model for our Universe.
When considering continuous global symmetries, the case that they are inconsistent can also be motivated by the black hole information paradox, in which case the argument can be generalized beyond the AdS/CFT correspondence [56]. One can loosely motivate this by arguing that continuous global symmetries imply conserved charges that are destroyed by black hole evaporation, therefore continuous global symmetries should not exist in nature, as we expect black hole evaporation to be unitary. This argument is more restrictive than the AdS/CFT result, as it applies to continuous global symmetries, such as U(1), but not global symmetries in general.
@@ II.3 Our Fiducial Model
We have four complex scalar fields A, B, C, and D, with four dark Higgs fields $\rm{H}_{1}$, $\rm{H}_{2}$, $\rm{H}_{3}$, and $\rm{H}_{4}$. We use complex scalar fields because charge assignments allow us to enforce a decay chain. One might initially suspect that a scalar field will not clump as dark matter because the kinetic terms could imply a sound speed of $c$, though there is no such problem [19], which we discuss further in the Appendix.
Since the theory has a U(1) $\times$ U(1) $\times$ U(1) $\times$ U(1) gauge group, we must introduce four gauge bosons to allow for gauge invariance: $\alpha_{\mu}$, $\beta_{\mu}$, $\gamma_{\mu}$, and $\delta_{\mu}$, where we define the Faraday tensors as $\Psi_{\mu\nu}=\partial_{\mu}\Psi_{\nu}-\partial_{\nu}\Psi_{\mu}$, and the covariant derivative as $\mathcal{D}_{\mu}\equiv\partial_{\mu}-i(g_{1}q_{1}\alpha_{\mu}+g_{2}q_{2}\beta_{\mu}+g_{3}q_{3}\gamma_{\mu}+g_{4}q_{4}\delta_{\mu})$. We note that a similar gauge group, U(1) $\times$ U(1) $\times$ U(1), can arise in string theory [57]. We also note that anomalies will not be present, as the model does not incorporate fermions [58]. We can express $\mathcal{L}$ as follows. For simplicity, we assume any mixed quartics of the form $(\mathrm{H}_{1}^{\dagger}\mathrm{H}_{1})(\mathrm{H}_{2}^{\dagger}\mathrm{H}_{2})$ have small coupling constants, and can be neglected.
$\mathcal{L}\supset$ $\;\mathcal{D}^{\mu}\mathrm{A}^{\dagger}\mathcal{D}_{\mu}\mathrm{A}-m_{\mathrm{A}}^{2}\,\mathrm{A}^{\dagger}\mathrm{A}+\mathcal{D}^{\mu}\mathrm{B}^{\dagger}\mathcal{D}_{\mu}\mathrm{B}-m_{\mathrm{B}}^{2}\,\mathrm{B}^{\dagger}\mathrm{B}$ (4)
$+\mathcal{D}^{\mu}\mathrm{C}^{\dagger}\mathcal{D}_{\mu}\mathrm{C}-m_{\mathrm{C}}^{2}\,\mathrm{C}^{\dagger}\mathrm{C}+\mathcal{D}^{\mu}\mathrm{D}^{\dagger}\mathcal{D}_{\mu}\mathrm{D}-m_{\mathrm{D}}^{2}\,\mathrm{D}^{\dagger}\mathrm{D}-\kappa\mathrm{H}_{1}^{\dagger}\mathrm{H}_{1}\sqrt{\mathrm{C}^{\dagger}\mathrm{C}}$
$-\Big(g_{\mathrm{AB}}\mathrm{H}_{2}^{\dagger}\mathrm{A}\mathrm{B}^{\dagger}\mathrm{B}+g_{\mathrm{BC}}\mathrm{H}_{3}^{\dagger}\mathrm{B}\mathrm{C}^{\dagger}\mathrm{C}+g_{\mathrm{CD}}\mathrm{H}_{4}^{\dagger}\mathrm{C}\mathrm{D}^{\dagger}\mathrm{D}+\text{h.c.}\Big)$
$-\sum_{\Psi_{\mu}\in\{\alpha_{\mu},\beta_{\mu},\gamma_{\mu},\delta_{\mu}\}}[\tfrac{1}{4}\!\Psi_{\mu\nu}\Psi^{\mu\nu}]+\sum_{i=1}^{4}\!\Big[\mathcal{D}^{\mu}\mathrm{H}_{i}^{\dagger}\mathcal{D}_{\mu}\mathrm{H}_{i}+m_{\mathrm{H}_{i}}^{2}\,\mathrm{H}_{i}^{\dagger}\mathrm{H}_{i}-\lambda_{\mathrm{H}_{i}}(\mathrm{H}_{i}^{\dagger}\mathrm{H}_{i})^{2}\Big]$
@@
Each time a U(1) is spontaneously broken, a dark Higgs field gains a non-zero vacuum expectation value. This then sequentially turns on the terms of the form $\sqrt{\mathrm{C}^{\dagger}\mathrm{C}}$, $\rm{A}\rm{B}^{\dagger}\mathrm{B}$, $\rm{B}\rm{C}^{\dagger}\mathrm{C}$, and $\rm{C}\rm{D}^{\dagger}\mathrm{D}$, where charge conservation forbids both two and three-body decays until the relevant U(1) symmetries are broken.
The terms $\rm{A}\rm{B}^{\dagger}\mathrm{B}$, $\rm{B}\rm{C}^{\dagger}\mathrm{C}$, and $\rm{C}\rm{D}^{\dagger}\mathrm{D}$ enable decays from one field to another, whereas the square root term allows C to have a negative equation of state parameter. The equation below yields the equation of state parameter $w(R)$ for a complex scalar field, which can be written in terms of its magnitude, $R$, and phase factor, $\phi$, giving $Re^{i\phi}$ [59].
@@
Here, the equation of state parameter is $w=0$ for a quadratic potential and w=1/3w=-1/3 for a potential that is linear in the magnitude, R, where the linear term dominates. One can alternatively obtain $0<w<1/3$ from semi-relativistic decay products. This is more natural than $w=1/3$ from a strongly-coupled quartic interaction that dominates the potential. If the field C is oscillating in a potential that has both a term linear in the magnitude (our square root term $\mathrm{H}_{1}^{\dagger}\mathrm{H}_{1}\sqrt{\mathrm{C}^{\dagger}\mathrm{C}}$) and a quadratic mass term, but where the former term dominates, one obtains $w\approx-1/3$. Since $\mathrm{H}_{1}$ obtains a VEV (vacuum expectation value) at early times, this square root term can naturally dominate after $\mathrm{U}_{1}$(1) is broken. We use this rather than a renormalizable term like $\mathrm{H}_{1}^{\dagger}\mathrm{C}$ because $\mathrm{H}_{1}^{\dagger}\mathrm{C}$ exhibits phase-dependence in addition to dependence on the magnitude, R, which complicates the dynamics. This square root term for a complex scalar is analogous to a potential $|\phi|$ for a real scalar $\phi$, as $R$ is strictly positive. Note that the square root term for C turns on before any decays begin, as the associated U(1) is broken first, though this term has no effect on the system until C is produced from decays later on.
@@ II.4 Self-Interacting Dark Matter Candidate From Our Fiducial Model
After the first U(1) is broken, yielding a dark Higgs with a mass of $\mathcal{O}$(MeV), the term $\mathcal{D}^{\mu}\rm{H_{1}}^{\dagger}\mathcal{D}_{\mu}\rm{H_{1}}$ produces an interaction of the form $gh\alpha_{\mu}\alpha^{\mu}$, which permits an $\alpha\alpha\to\alpha\alpha$ scattering processes via a scalar mediator, as shown below. Note that $h$ is the radial mode of the dark Higgs, after symmetry breaking gives it a mass. For $\alpha\sim\rm{GeV}$ and $h\sim\rm{MeV}$, our fiducial model produces a natural SIDM candidate, one that could conceivably address some of the issues with CDM if $\alpha$ comprises the bulk of the dark matter. We show the $t$ and $u$-channel Feynman diagrams in Figure 1.
@@ III.1 H(z) From Our Fiducial Model
We have equation of state parameters $w_{\rm{A}}=0$, $w_{\rm{B}}=\frac{1}{12}$, $w_{\rm{C}}=-\frac{1}{3}$, and $w_{\rm{D}}=\frac{1}{12}$, where B and D are warm dark matter. We use $\Gamma_{\rm{A}}=\Gamma_{\rm{B}}=\Gamma_{\rm{C}}=10^{-15}$ $\rm{sec^{-1}}$. (See the Appendix for the equation that expresses the decay rate in terms of the masses and coupling constants.) In the equations below, $\rho_{\mathrm{A}}$, $\rho_{\mathrm{B}}$, $\rho_{\mathrm{C}}$, and $\rho_{\mathrm{D}}$ are the densities of the complex scalars in the cascade, whereas $\rho_{\mathrm{O}}$ is the density of the other matter (baryons+other dark matter), $\rho_{r}$ is the radiation density, and $\rho_{\Lambda}$ is the dark energy density. In the equations below, $t_{2}$, $t_{3}$, and $t_{4}$ correspond to redshifts of 3, 1.5, and 0.6. We note that the square root term for C (which gives it $w<0$) turns on at early times when $\alpha_{\mu}$ gains a mass, though this term does not have an effect until C is produced from decays of B. We neglect any energy transfer to the massive radial modes of the dark Higgs fields via three-body decays after the relevant U(1) symmetries are broken; a more detailed analysis could consider this.
$\dot{\rho}_{\rm A}+3H(1\!+\!w_{\rm A})\rho_{\rm A}$ $=-\Theta(t\!-\!t_{2})\Gamma_{\rm A}\rho_{\rm A}$ (6a)
$\dot{\rho}_{\rm B}+3H(1\!+\!w_{\rm B})\rho_{\rm B}$ $=\Theta(t\!-\!t_{2})\Gamma_{\rm A}\rho_{\rm A}$
$\quad-\Theta(t\!-\!t_{3})\Gamma_{\rm B}\rho_{\rm B}$ (6b)
$\dot{\rho}_{\rm C}+3H(1\!+\!w_{\rm C})\rho_{\rm C}$ $=\Theta(t\!-\!t_{3})\Gamma_{\rm B}\rho_{\rm B}$
$\quad-\Theta(t\!-\!t_{4})\Gamma_{\rm C}\rho_{\rm C}$ (6c)
$\dot{\rho}_{\rm D}+3H(1\!+\!w_{\rm D})\rho_{\rm D}$ $=\Theta(t\!-\!t_{4})\Gamma_{\rm C}\rho_{\rm C}$ (6d)
@@
We integrate the system of equations from recombination at $z_{\rm rec}=1090$ to $z=0$, where we specify our initial conditions by setting $h_{i}$ = 0.674, where $\rm{H_{0}}$ is $100\rm{\frac{km}{secMpc}}$$h_{i}$ in units of $\rm{sec}^{-1}$. The radiation $\Omega_{r}$ and matter $\Omega_{m}$ components are dictated by the equations below [60, 61], where $\Omega_{\Lambda}$ follows from flatness:
@@
$\rho_{\Lambda,i}=\Omega_{\Lambda}\rho_{\rm crit,0}$ (14)
III.2 Dynamical Dark Energy
We use the equation below to calculate H(z) for a Universe with cold dark matter and dynamical dark energy [62]. We plot H(z) from our model alongside $w_{0}w_{a}$CDM, which has the equation of state parameter $w_{\rm{DE}}(a)=w_{0}+w_{a}(1-a)$ for dark energy, to show that our model could display qualitative consistency with the DESI DR2 trends. In a following work, we plan to replace the cosmological constant with a quintessence field (w>1w>-1) that is not coupled to the dark matter. This is appealing for avoiding the formation of a cosmological horizon, though it moves beyond a realization of Chen and Loeb’s ideas, so we leave it to a separate paper. Similar to Chen and Loeb [19], who deferred a fit to DESI data to a separate study [20], we leave a full parameter fit to DESI results for future work. Such a fit is beyond the scope of the current study to explore the space of potentially viable models.